OpenStudy (anonymous):

Given the trinomial, what is the value of the coefficient 'B' in the factored form? 2x2 - 12xy - 32y2 = 2(x - 8y)(x + By) B= ____________ I am not looking for an answer, I already know it. I just need to understand how to complete the problem.

OpenStudy (anonymous):

2x^2-12xy-32y^2=2(x-8y)(x+By)

OpenStudy (anonymous):

@alyssa_xo

OpenStudy (bibby):

as luck would have it, I'm alyssa_xo

OpenStudy (anonymous):

LOL! >.> you are never online you must help me!

OpenStudy (bibby):

cause this is my main account lol anyhow you're given the factorization on the right, use that

OpenStudy (bibby):

for example you know you can pull out a 2 from each term

OpenStudy (anonymous):

i dont understand. I can do that on the left, but the right? I foil first? right?

OpenStudy (anonymous):

correct...not right... thats confusing

OpenStudy (bibby):

it's a statement of equivalence. they're the same thing

OpenStudy (bibby):

so if you multiply out the right you get the thing on the left

OpenStudy (bibby):

assuming you know what B is

OpenStudy (anonymous):

yea, on the left i get 2(x^2-6xy-16y^2)

OpenStudy (bibby):

can you factor the polynomial thing

OpenStudy (anonymous):

the right i foil, but the 2 is already taken out, like on the left

OpenStudy (bibby):

no, don't touch the right, it's a blueprint

OpenStudy (anonymous):

kay so leave it as 2(x - 8y)(x + By)

OpenStudy (anonymous):

so now i have 2(x^2-6xy-16y^2)= 2(x - 8y)(x + By) What do i do next?

OpenStudy (bibby):

the goal is to factor the left until it resembles the form presented on the right

OpenStudy (anonymous):

wait but there is xy and y^2 i dont know how to factor that sorry

OpenStudy (bibby):

I'd tell you to factor x^2-6xy-16y^2 but I can't remember how to factor that so just divide

OpenStudy (bibby):

\(\dfrac{ x^2-6xy-16y^2}{x-8y}\)

OpenStudy (anonymous):

oh okay sort of like 2=2x and solving for x

OpenStudy (anonymous):

uhm.... i got x-6-2y

OpenStudy (anonymous):

sorry x-6+2y

OpenStudy (bibby):

x+2y

OpenStudy (anonymous):

what about the 6? how did it disappear?

OpenStudy (bibby):

where'd you get the -6 from

OpenStudy (bibby):

I can't finish it as I have to play a game

OpenStudy (bibby):

can u show ur work on the division